OpenStudy (anonymous):

PLEASE ANYONE WHO CAN HELP ME!! algebra 2: 2r- 11/5 <= 5/4r-11/4 or 47/10r+29/8 >-2r +9/8

OpenStudy (anonymous):

What are you trying to do?

OpenStudy (anonymous):

I'M TRYING TO SOLVE THE INEQUALITY AND GRAPH ITS SOLUTION

OpenStudy (anonymous):

You can put it into y = mx+b form to graph it (equation of a straight line). Solve for r.

OpenStudy (anonymous):

I'm only doing compound inequality then graphing it on a number line

OpenStudy (anonymous):

can u still help?

OpenStudy (anonymous):

Then solve for r, and graph it on a number line, but show your work if you want further help.

OpenStudy (anonymous):

i know you have to solve for r I'm kinda asking how you would start solving i guess like what would be the first step to solving

OpenStudy (anonymous):

Is your question, \[2r-\frac{ 11 }{ 5 } \le \frac{ 5 }{ 4 }3-\frac{ 11 }{ 4 }\]

OpenStudy (anonymous):

that should be a r not a 3

OpenStudy (anonymous):

Collect like terms, it's just like a regular problem with a different sign.